# LeanKAN: A Parameter-Lean Kolmogorov-Arnold Network Layer with Improved Memory Efficiency and Convergence Behavior

### v1

@@ IV.1.3 Higher Order Multiplication
Such benefits become more significant when a higher order of multiplication is needed. For example, in the same problem setup, an additional output node $z_{5}=x_{1}x_{2}x_{3}$ could be learned simply by selecting the structural hyperparameter $n^{mu}=3$ in our LeanKAN, which does not change the number of activation functions or parameters. Interpretation-wise, we might expect to see three zero-offset identity functions in the fifth nodes for these three inputs. For MultKAN, however, $k$ would need to increase to 3 to accommodate third-order multiplication, resulting in 16 additional activation functions compared to LeanKAN (and $2$ times the total activation parameters). As a result, a single MultKAN layer risks blowing out the number of activation functions and dummy activations to learn higher order multiplication, worsening the memory efficiency and training efficiency while also concealing the interpretation of the true underlying behavior under a series of extraneous activations and operations. Partial mitigation of this issue is perhaps possible by optimizing the two hyperparameters, $k$ and $1-n^{a}$, for the order of multiplication and the number of output multiplication nodes, respectively. That being said, these two degrees of freedom add significant complexity to the network designers’ hyperparamer tuning task, and as discussed above $k=3$ is already the most parameter-lean value possible for third-order multiplication. Conversely, LeanKAN has only one hyperparameter $n^{mu}$ for the number of output multiplication nodes, with full utilization of all multiplication inputs (i.e. the order of multiplication is equal to the number of multiplication inputs), and with no changes to the parameter counts regardless of the value of $n^{mu}$.
@@
For practical use, we suggest setting $n^{mu}$ to half of the input size initially, for a balance between addition and multiplication nodes. Physical insights in the specific problem at hand may lead users to increase or decrease this value, at no parameter cost. Additionally, following from the discussion above, we suggest general use of LeanKAN in a two-layer structure: an input AddKAN layer to compose hidden node values that distribute all inputs across the entire width of the hidden layer, and then an output LeanKAN layer to inject sparse multiplication into the input-diverse hidden layer values. In doing so, users can completely negate the input-splitting issue seen in $z_{2}$ and $z_{4}$ in Figs. 2(B) and 2(D). In an extreme case where physical insights support the sole use of multiplication, a single-layer LeanKAN with $n^{mu}=n$ can also avoid this issue by applying multiplication to all nodes (similarly to how an AddKAN avoids this issue by applying addition to all nodes). By taking one of these two approaches, users can benefit from the substantial parameter savings of LeanKAN at a change in accuracy and training dynamics that, according to the results of Fig. 3, appears to be favorable to LeanKAN. In the next section, we learn simple biological dynamics via LeanKAN in an augmented architecture to invesigate whether its lean parameterization can scale up to a more realistic case and continue to provide these counterintuitive performance improvements.
IV.2 LeanKAN in Orindary Differential Equations
@@
In this section, we further investigate the quantitatively improved performance of LeanKAN compared to MultKAN in the respective optimal configurations: MultKAN is studied as the first of two layers (with the second being a standard additive KAN), and LeanKAN is studied as the second of two layers (the first similarly being a standard additive KAN). In doing so, all issues regarding input or output multiplication limitations (as studied in Fig. 2) are resolved.
@@
In this case study, we look into the classical Lotka-Volterra predator-prey model using the Kolmogorov-Arnold Network Ordinary Differential Equation (KAN-ODE) framework of Koenig et al. [5]. Our aim here is to demonstrate LeanKAN as an effective layer structure in augmented KAN-based architectures (like KAN-ODEs). Briefly, and as was presented in more detail in Koenig et al. [5], the governing equations given the state variables $\mathbf{u}=[x,y]$ are
@@
with $\alpha=1.5$, $\beta=1$, $\gamma=1$, and $\delta=3$ and an initial condition of $\mathbf{u}_{0}=[x_{0},y_{0}]=[1,1]$. The training data is generated in time span of $t\in[0,3.5]$ s, while the unseen testing data is generated in the time span of $t\in[3.5,14]$ s. The KAN-ODEs learn the underlying model as
@@
where $\bm{\theta}$ is the collection of all trainable parameters in the KAN. Using the mean squared error (MSE) loss function and the adjoint sensitivity analysis [19, 20] as in Koenig et al. [5], we update $\bm{\theta}$ from a random initialization to fit the training data. The ADAM optimizer [18] was adopted with a learning rate of $2\cdot 10^{-4}$ for both LeanKAN-based and MultKAN-based architectures in the fully converged cases of Secs. IV.2.1 and IV.2.2, and $5\cdot 10^{-3}$ in the rapid convergence test of Sec. IV.2.3.
IV.2.1 Training Improvements with Large KANs
@@
Figure 4: Fully converged Lotka-Volterra inference comparisons between LeanKAN and MultKAN. (A) KAN-ODE layer structures referenced in the remaining subfigures. AddKAN is reported fully in [5]. (B-C) Training and testing results, respectively, with different traces representing the MultKAN and LeanKAN profiles. Both structures have a single hidden layer of 10 nodes, half of which are multiplication nodes, on a 5-point grid (300 MultKAN parameters and 240 LeanKAN parameters). (D-E) Neural convergence study comparing the two KAN layer architectures at various node and grid layouts, with the AddKAN convergence results of [5] plotted in grey. LeanKAN is seen to outperform MultKAN at all tested sizes, despite having $25\%$ fewer parameters for a given size. Linear fit to convergence results suggest that while convergence rates are similar, LeanKAN has fewer extraneous parameters, shifting its linear fit significantly toward smaller parameter counts (see orange arrows for the average quantitative impact on the number of parameters and loss metrics). Linear fits for LeanKAN and AddKAN include only those points occurring before loss saturation.
@@
We first note the optimal KAN structure reported in Koenig et al. [5], which had a two-node input, a ten-node hidden layer, and a two-node output layer, all using standard AddKAN connections. All activations had five grid points each, and as in the first example here used RBF basis functions with a Swish basis residual function. This KAN contained 240 parameters. Here, we use it as the basis on which to design MultKAN and LeanKAN based networks.
@@
As proposed in Sec. IV.1, the LeanKAN model used here involves replacing the second layer of the original KAN structure of [5]. Instead of 20 additive KAN activations connecting the ten hidden nodes with the two output nodes, we choose $n^{mu}=5$ to split the input nodes half and half between additive and multiplicative KAN operations. The overall structure remains as 2 nodes into 10 nodes into 2 nodes, for a conserved total of 240 parameters.
@@
Separately, the MultKAN model used here involves replacing the first (input) layer of the original KAN structure of [5]. We similarly choose $n^{a}=5$ to split the hidden nodes half and half between identity and multiplication operations. With the smallest multiplication order possible in the MultKAN framework, $k=2$, the MultKAN parameter inflation leads to 300 total parameters, rather than 240 (half of the activations in the formerly 120-parameter MultKAN layer are doubled, for a total increase of 60 parameters). We also remark that $k=2$, in a certain sense, contains inductive bias relevant to the problem at hand. Knowing that the governing equations in Eq. 16 contain second-order multiplication only, the MultKAN is encoded to contain only second-order multiplication structures. LeanKAN, on the other hand, has an order of multiplication equal to $n^{mu}$, and does not benefit from a problem-knowledge-enforced multiplication order. Both two-layer structures (as well as the original AddKAN structure of [5]) are visualized in Fig. 4(A).
@@
Both models are trained for a relatively long $10^{6}$ epochs, which in both cases takes on the order of tens of hours on a single CPU. Mean squared error results are compared across these network structures in Figs. 4(B-C) for training and testing losses, respectively. In the early epochs of training (up to and around $2\cdot 10^{5}$ epochs), we see a remarkable improvement in training loss behavior from the MultKAN to the LeanKAN. Both converge fairly well to low loss, but the LeanKAN is rapidly able to reach a minimum value two orders of magnitude smaller than that of the MultKAN in these early epochs. In the testing loss results of Fig. 4(C), we see an even more pronounced discrepancy in the extremely early epochs, where the LeanKAN is almost immediately able to jump down to the $10^{-4}$ MSE range, while the MultKAN experiences a rather long delay while it finds its way below even just an MSE value of $1$.
Looking out to the rest of the training window, we see that MultKAN is able to break into the $10^{-5}$ MSE range on training error after around $5\cdot 10^{5}$ epochs, but is still never able to reach the sub-$10^{-6}$ value that the LeanKAN reaches in just $2\cdot 10^{5}$ epochs. The testing losses, once the MultKAN is given enough time to converge, are fairly similar. It should be highlighted that the LeanKAN appears to have matched or outperformed the MultKAN with 60 fewer parameters (240 for LeanKAN and 300 for MultKAN).
@@
To summarize, in this case with a relatively large KAN size and a long training window, we find that LeanKAN appears to outperform MultKAN in most aspects: it achieves a lower training loss, converges faster, and has a lower parameter count and complexity (even with the minimum $k=2$ for the MultKAN). In the two following sections, we further probe these noteworthy performance improvements. First, to further elucidate the increased converged performance seen in the LeanKAN and test its robustness, we carry out a traditional neural convergence study to evaluate this behavior across networks of various node counts and grid densities. Then, to further study the improved early-epoch training dynamics seen in Figs. 4(B-C), we carry out a heavily restricted training cycle with smaller networks and just 7,000 epochs. We alternate in these two comparisons between identical network structures (with MultKAN benefitting from more parameters) and identical parameter counts (with MultKAN facing reduced nodes due to its extraneous activations) to further probe the parameter size discrepancy between architectures.
IV.2.2 Neural Convergence Comparison
@@
We have identified in the results of Sec. IV.1 as well as Sec. IV.2.1 that even with reduced parameter counts, LeanKAN appears to be able to slightly outperform MultKAN. We believe that this performance boost stems largely from the improved parameter efficiency of LeanKAN. While LeanKAN and MultKAN should both be able to converge to similar levels of loss when given arbitrarily many nodes and parameters, it is noteworthy that LeanKAN can do so with what appears to be significantly fewer parameters thanks to its elimination of dummy activations and extraneous parameters. To further test this behavior, we show neural convergence results for this example in Figs. 4(D-E). Testing across a range of identical node and grid structures of varying parameter counts, we see nearly identical neural scaling orders between MultKAN and LeanKAN (5.2 vs. 5.3, respectively). However, the entire curve for LeanKAN is shifted significantly to the left, indicating that while both are capable of achieving the same neural scaling and final strong performance, LeanKAN is simply able to do so with fewer parameters. In fact, the LeanKAN was able to converge to its minimum loss value (where larger networks provide no additional accuracy) with just 168 parameters, while the MultKAN appears to be not yet fully converged even with 300 parameters. Specifically, for a given MultKAN parameter count and converged loss, LeanKAN is capable of achieving the same converged loss with $2.7$ times fewer parameters on average. Alternatively, the LeanKAN with the same number of parameters is capable of achieving $200$ times less converged loss on average.
@@
Figure 5: Transient training dynamics comparison between LeanKAN and MultKAN for Lotka-Volterra system. (A-B) Training and testing results, respectively, with small KAN sizes (100 parameters) and 7000 epochs of training only. In addition to improved converged results (Fig. 4), LeanKAN is seen to have further improved transient training dynamics. (C) Training and testing data reconstruction using the results of (A-B). Both KAN architectures appear to predict the training data sufficiently well, though MultKAN strugges more with the unseen testing data.
IV.2.3 Increased Convergence Speed
@@
The last behavior that we noted in Figs. 4(B-C) is the improved transient training dynamics of LeanKAN. To further test this behavior, we replace the large-model, long-window training used in Figs. 4(B-C) with smaller models and shorter training times that are more indicative of potential real-world application (especially in a case with simple 1-D trajectories like is studied here).
@@
In all previous cases, we used identical node structures in MultKAN and LeanKAN, which inherently leads to larger MultKAN parameter counts. Here, we instead equalize the number of parameters in both networks. The MultKAN structure is reduced to four hidden nodes (with $n^{a}$=2 for two multiplication and two addition nodes) and four gridpoints, with 100 total parameters. The LeanKAN structure, meanwhile, is reduced to five hidden nodes (with $n^{mu}=3$ for three multiplication and two addition nodes) with the same four gridpoints, resulting in 100 total parameters. This reduction of LeanKAN to the same parameter count with an additional hidden node compared to MultKAN is possible thanks to its compact form compared to even the smallest ($k=2$) MultKAN, and its ability to represent multiplication nodes using fewer parameters and fewer activations.
@@
Both networks in this case are trained for relatively small 7,000 epochs (chosen to allow the MultKAN enough training time to fit the testing data passably well). Training dynamics are shown in Figs. 5(A-B). Looking first at the training loss and training data reconstruction, we notice that the LeanKAN training loss rapidly drops below that of the MultKAN and is nearly an order of magnitude less than that of the MultKAN for the entire training window. This corroborates the behavior seen at early epochs in Fig. 4(B). In the data reconstruction of Fig. 5(C), both fit the training data strongly on inspection, although a very close look shows that the MultKAN, especially toward the end of the 3.5 s training window, begins to diverge from the data.
@@
Looking next at the testing data, we see further corroboration of the distinction originally noted in Fig. 4(C). That is, the LeanKAN is rapidly able to capture the behavior of the underlying model and drop its testing loss to acceptably low values, while the MultKAN struggles and does not begin to notably decrease its testing loss until after 5,000 epochs have passed. In the testing window reconstruction of Fig. 5(C), this distinction is obvious. The LeanKAN fits the unseen data very well and does not begin to deviate notably until the last oscillation, while the MultKAN almost immediately diverges from the testing data and has significant error by the final oscillation.
@@
To summarize, the three tests carried out using the Lotka-Volterra equations in the KAN-ODE framework found that LeanKAN appears to outperform MultKAN in most metrics across the board. Both are able to achieve strong accuracy in most cases, but LeanKAN is able to converge faster, to lower loss values, and with significantly fewer parameters than MultKAN. Some of the bulk in MultKAN that leads to these performance drawbacks is likely able to be pruned downstream, as studied in [1, 2, 5], although in the cases studied here we saw no reason to begin optimization at such a heavily parameterized structure in the first place, as the leaner and still-prunable LeanKAN saw superior performance across all metrics.
V Conclusions

### v2

@@ 4.1.3 Higher Order Multiplication
Such benefits become more significant when a higher order of multiplication is needed. For example, in the same problem setup, an additional output node $z_{5}=x_{1}x_{2}x_{3}$ could be learned simply by selecting the structural hyperparameter $n^{\text{mu}}=3$ in LeanKAN, which does not change the number of activation functions or parameters. Interpretation-wise, we might expect to see three zero-offset identity functions in the fifth node for these three inputs. For MultKAN, however, $k$ would need to increase to 3 to accommodate third-order multiplication, resulting in 16 additional activation functions compared to LeanKAN (and $2$ times the total activation parameters). As a result, hyperparameter tuning of $k$ in MultKAN layers is not only complex due to the competing tuning of the $1-n^{a}$ hyperparameter, but also due to the tradeoff between multiplication order and memory efficiency, as higher-order multiplication risks blowing out the number of extraneous or dummy activations. LeanKAN, meanwhile, has a single hyperparameter $n^{\text{mu}}$ that has no impact on parameter count.
@@
For practical use, we suggest setting $n^{\text{mu}}$ to half of the input size initially, for a balance between addition and multiplication nodes. Physical insights in the specific problem at hand may lead users to increase or decrease this value, at no parameter cost. For wider layers with many nodes, reducing $n^{\text{mu}}$ may help to accelerate training by reducing the order of multiplication. Additionally, following from the discussion above, we suggest general use of LeanKAN in a two-layer structure: an input AddKAN layer to compose hidden node values that distribute all inputs across the entire width of the hidden layer, and then an output LeanKAN layer to inject sparse multiplication into the input-diverse hidden layer values. In doing so, users can completely negate the input-splitting issue seen in $z_{2}$ and $z_{4}$ in Figs. 2(B) and 2(D). In an extreme case where physical insights support the sole use of multiplication, a single-layer LeanKAN with $n^{\text{mu}}=n$ can also avoid this issue by applying multiplication to all nodes (similarly to how an AddKAN avoids this issue by applying addition to all nodes). By taking one of these two approaches, users can benefit from the substantial parameter savings of LeanKAN at a change in accuracy and training dynamics that, according to the results of Fig. 3, appears to be favorable to LeanKAN. We finally remark that LeanKAN and MultKAN are not mutually exclusive. While in the remainder of this manuscript we use the two separately in order to facilitate comparison, it is possible to design a two-layer network where a MultKAN layer feeds into a LeanKAN layer, providing addition and multiplication at both layers across all nodes. We do not currently believe this hybrid structure to be promising due to the extraneous bulk in MultKAN, although we describe it here for completeness. In the next section, we learn simple biological dynamics via LeanKAN in an augmented architecture to investigate whether its lean parameterization can scale up to a more realistic case and continue to provide these somewhat counterintuitive performance improvements.
4.2 LeanKAN in Ordinary Differential Equations
@@
In this section, we further investigate the quantitatively improved performance of LeanKAN compared to MultKAN in the respective optimal configurations: MultKAN is studied as the first of two layers (with the second being an AddKAN layer), and LeanKAN is studied as the second of two layers (the first similarly being an AddKAN layer). In doing so, all issues regarding input or output multiplication limitations (as studied in Fig. 2) are resolved.
@@
In this case study, we look into the classical Lotka-Volterra predator-prey model using the Kolmogorov-Arnold Network Ordinary Differential Equation (KAN-ODE) framework [13]. Our aim here is to demonstrate LeanKAN as an effective layer structure in augmented KAN-based architectures (such as KAN-ODEs). Briefly, and as was presented in more detail in [13], the governing equations given the state variables $\mathbf{u}=[x,y]$ are
@@
with $\alpha=1.5$, $\beta=1$, $\gamma=1$, and $\delta=3$ and an initial condition of $\mathbf{u}_{0}=[x_{0},y_{0}]=[1,1]$. The training data is generated in time span of $t\in[0,3.5]$ s, while the unseen testing data is generated in the time span of $t\in[3.5,14]$ s. All data was generated using the Tsitouras 5/4 Runge-Kutta method [27] and a step size of 0.1 s. The KAN-ODEs learn the underlying model as
@@
where $\bm{\theta}$ is the collection of all trainable parameters in the KAN. Using the mean squared error (MSE) loss function and the adjoint sensitivity analysis [4, 23] as in [13], we update $\bm{\theta}$ from a random initialization to fit the training data. The ADAM optimizer [12] was adopted with a learning rate of $2\times 10^{-4}$ for both LeanKAN-based and MultKAN-based architectures in the fully converged cases of Secs. 4.2.1 and 4.2.2, and $5\times 10^{-3}$ in the rapid convergence test of Sec. 4.2.3.
4.2.1 Training Improvements with Large KANs
@@
Figure 4: Fully converged Lotka-Volterra inference comparisons between LeanKAN and MultKAN. (A) KAN-ODE layer structures referenced in the remaining subfigures. AddKAN is reported fully in [13]. (B-C) Training and testing results, respectively, with different traces representing the MultKAN and LeanKAN profiles. Both structures have a single hidden layer of 10 nodes, half of which are multiplication nodes, on a 5-point grid (300 MultKAN parameters and 240 LeanKAN parameters). (D) Neural convergence study comparing the two KAN layer architectures at various node and grid layouts, with the AddKAN convergence results of [13] plotted in grey. LeanKAN is seen to outperform MultKAN at all tested sizes, despite having $25\%$ fewer parameters for a given size. Linear fit to convergence results suggest that while convergence rates are similar, LeanKAN has fewer extraneous parameters, shifting its linear fit significantly toward smaller parameter counts (see orange arrows for the average quantitative impact on the number of parameters and loss metrics). Linear fits for LeanKAN and AddKAN include only those points occurring before loss saturation.
@@
We first note the optimal KAN structure reported originally [13], which had a two-node input, a ten-node hidden layer, and a two-node output layer, all using standard AddKAN connections. All activations had five grid points each, and as in the first example here used RBF basis functions with a Swish basis residual function. This KAN contained 240 parameters. Here, we use it as the basis on which to design MultKAN and LeanKAN based networks.
@@
As proposed in Sec. 4.1, the LeanKAN model used here involves replacing the second layer of the original KAN structure of [13]. Instead of 20 additive KAN activations connecting the ten hidden nodes with the two output nodes, we choose $n^{\text{mu}}=5$ to split the input nodes half and half between additive and multiplicative KAN operations. The overall structure remains as 2 nodes into 10 nodes into 2 nodes, for a conserved total of 240 parameters.
@@
Separately, the MultKAN model used here involves replacing the first (input) layer of the original KAN structure of [13]. We similarly choose $n^{a}=5$ to split the hidden nodes half and half between identity and multiplication operations. With the smallest multiplication order possible in the MultKAN framework, $k=2$, the MultKAN parameter inflation leads to 300 total parameters, rather than 240 (half of the activations in the formerly 120-parameter MultKAN layer are doubled, for a total increase of 60 parameters). We also remark that $k=2$, in a certain sense, contains inductive bias relevant to the problem at hand. Knowing that the governing equations in Eq. 18 contain second-order multiplication only, the MultKAN is encoded to contain only second-order multiplication structures. LeanKAN, on the other hand, has an order of multiplication equal to $n^{\text{mu}}$=5, and does not benefit from a problem-knowledge-enforced multiplication order. Both two-layer structures as well as the original AddKAN structure of [13] are visualized in Fig. 4(A).
Table 2: Neural convergence results.
Structure MultKAN LeanKAN (ours)
Nodes $\bm{n^{\text{mu}}}$ or $\bm{n^{m}}$ Grid Param. count MSE Param. count MSE
4 2 3 80 1.3×$\times 10^{-3}$ 64 1.5×$\times 10^{-5}$
5 3 5 156 2.7×$\times 10^{-6}$ 120 2.8×$\times 10^{-7}$
6 3 5 210 6.7×$\times 10^{-6}$ 168 1.0×$\times 10^{-7}$
10 5 5 300 1.0×$\times 10^{-6}$ 240 3.1×$\times 10^{-7}$
@@
Both models are trained for $2\times 10^{5}$ epochs, a relatively large number which in both cases takes on the order of tens of hours on a single CPU. Mean squared error results are compared across these network structures in Figs. 4(B-C) for training and testing losses, respectively. There, a remarkable improvement in training loss behavior is observed when moving from the MultKAN to the LeanKAN. Both converge fairly well to low loss, but the LeanKAN is rapidly able to reach a minimum value two orders of magnitude smaller than that of the MultKAN. In the testing loss results of Fig. 4(C), we see an even more pronounced discrepancy in the extremely early epochs, where the LeanKAN is almost immediately able to jump down to the $10^{-4}$ MSE range, while the MultKAN experiences a rather long delay while it finds its way below even just an MSE value of $1$. While the final converged testing losses are not as dramatically different as the training losses, LeanKAN retains an advantage of nearly a full order of magnitude.
@@
To summarize, in this case with a relatively large KAN size and a long training window, we find that LeanKAN appears to outperform MultKAN in most aspects, even with 60 fewer parameters: it achieves a lower training loss, converges faster, and has a lower parameter count and complexity (even with the minimum $k=2$ for the MultKAN). In the two following sections, we further probe these noteworthy performance improvements. First, to further elucidate the increased converged performance seen in the LeanKAN and test its robustness, we carry out a traditional neural convergence study to evaluate this behavior across networks of various node counts and grid densities. Then, to further study the improved early-epoch training dynamics seen in Figs. 4(B-C), we carry out a heavily restricted training cycle with smaller networks and just 7,000 epochs. We alternate in these two comparisons between identical network structures (with MultKAN benefitting from more parameters) and identical parameter counts (with MultKAN facing reduced nodes due to its extraneous activations) to further probe the parameter size discrepancy between architectures.
4.2.2 Neural Convergence Comparison
@@
We have identified in the results of Sec. 4.1 as well as Sec. 4.2.1 that even with reduced parameter counts, LeanKAN appears to be able to slightly outperform MultKAN. We believe that this performance boost stems largely from the improved parameter efficiency of LeanKAN. While LeanKAN and MultKAN should both be able to converge to similar levels of loss when given arbitrarily many nodes and parameters, it is noteworthy that LeanKAN can do so with what appears to be significantly fewer parameters thanks to its elimination of dummy activations and extraneous parameters. To further test this behavior, we show neural convergence results for this example in Fig. 4(D), with numerical values shown in Table 2. Testing across a range of identical node and grid structures of varying parameter counts, we see nearly identical neural scaling orders between MultKAN and LeanKAN (5.2 vs. 5.3, respectively). However, the entire curve for LeanKAN is shifted significantly to the left, indicating that while both are capable of achieving the same neural scaling and final strong performance, LeanKAN is simply able to do so with fewer parameters. In fact, the LeanKAN was able to converge to its minimum loss value (where larger networks provide no additional accuracy) with just 168 parameters, while the MultKAN appears to be not yet fully converged even with 300 parameters. Specifically, for a given MultKAN parameter count and converged loss, LeanKAN is capable of achieving the same converged loss with $2.7$ times fewer parameters on average. Alternatively, the LeanKAN with the same number of parameters is capable of achieving $200$ times less converged loss on average.
@@
Figure 5: Transient training dynamics comparison between LeanKAN and MultKAN for Lotka-Volterra system. (A-B) Training and testing results, respectively, with small KAN sizes (100 parameters) and 7000 epochs of training only. In addition to improved converged results (Fig. 4), LeanKAN is seen to have further improved transient training dynamics. (C) Training and testing data reconstruction using the results of (A-B). Both KAN architectures appear to predict the training data sufficiently well, though MultKAN struggles with the unseen testing data.
4.2.3 Increased Convergence Speed
@@
The last behavior that we noted in Figs. 4(B-C) is the improved transient training dynamics of LeanKAN. To further test this behavior, we replace the large-model, long-window training used in Figs. 4(B-C) with smaller models and shorter training times that are more indicative of potential real-world application (especially in a case with simple 1-D trajectories as is studied here).
@@
In all previous cases, we used identical node structures in MultKAN and LeanKAN, which inherently leads to larger MultKAN parameter counts. Here, we instead equalize the number of parameters in both networks. The MultKAN structure is reduced to four hidden nodes (with $n^{a}$=2 for two multiplication and two addition nodes) and four gridpoints, with 100 total parameters. The LeanKAN structure, meanwhile, is reduced to five hidden nodes (with $n^{\text{mu}}=3$ for three multiplication and two addition nodes) with the same four gridpoints, resulting in 100 total parameters. This reduction of LeanKAN to the same parameter count with an additional hidden node compared to MultKAN is possible thanks to its compact form compared to even the smallest ($k=2$) MultKAN, and its ability to represent multiplication nodes using fewer parameters and fewer activations.
@@
Both networks in this case are trained for a relatively shorter period of 7,000 epochs (chosen to allow the MultKAN enough training time to fit the testing data passably well). Training dynamics are shown in Figs. 5(A-B), with reconstructed profiles in Fig. 5(C). Looking first at the training loss and training data reconstruction, we notice that the LeanKAN training loss rapidly drops below that of the MultKAN and is nearly an order of magnitude less than that of the MultKAN for the entire training window. This corroborates the behavior seen at early epochs in Fig. 4(B). In the data reconstruction of Fig. 5(C), both fit the training data strongly on inspection, although a very close look shows that the MultKAN, especially toward the end of the 3.5 s training window, begins to diverge from the data.
@@
Looking next at the testing data, we see further corroboration of the distinction originally noted in Fig. 4(C). That is, the LeanKAN is rapidly able to capture the behavior of the underlying model and drop its testing loss to acceptably low values, while the MultKAN struggles and does not begin to notably decrease its testing loss until after 5,000 epochs have passed. In the testing window reconstruction of Fig. 5(C), this distinction is obvious. The LeanKAN fits the unseen data very well and does not begin to deviate notably until the final oscillation, while the MultKAN almost immediately diverges from the testing data and has significant error by the final oscillation.
@@
To summarize, the three tests carried out using the Lotka-Volterra equations in the KAN-ODE framework found that LeanKAN appears to outperform MultKAN in most metrics across the board. Both are able to achieve strong accuracy in most cases, but LeanKAN is able to converge faster, to lower loss values, and with significantly fewer parameters than MultKAN. Some of the bulk in MultKAN that leads to these performance drawbacks is likely able to be pruned downstream, as studied in [17, 16, 13], although in the cases studied here we saw no reason to begin optimization at such a heavily parameterized structure in the first place, as the leaner and still-prunable LeanKAN saw superior performance across all metrics. For further LeanKAN evaluation, we provide a Lotka-Volterra learning case with noisy and irregular data in the Appendix.
4.3 LeanKAN in Partial Differential Equations (PDEs)
We conclude with a brief demonstration of LeanKAN’s capability for PDE learning, adapting the complex-valued Schrödinger problem used in quantum mechanics and previously studied with AddKAN layers [13],
${i\frac{\partial u}{\partial t}+\frac{1}{2}\frac{\partial^{2}u}{\partial x^{2}}+\lvert u\rvert^{2}u=0},$ (20)
where $u\in\mathbb{C}^{1}$. The time and spatial domains used are $x\in[-5,5]$ and $t\in[0,\pi/2]$. The initial condition and boundary conditions are defined as
${u\left(x,0\right)}$ $={\text{sech}\left(x\right)},$ (21)
${u\left(-5,t\right)}$ $={u\left(5,t\right)},$ (22)
${\frac{\partial u}{\partial x}\left(-5,t\right)}$ $={\frac{\partial u}{\partial x}\left(5,t\right)}.$
Figure 6: Schrödinger Equation case study. (A) True solution, LeanKAN prediction, LeanKAN prediction error, and AddKAN prediction error (top to bottom), for the real component, imaginary component, and magnitude of the solution (left to right). AddKAN and LeanKAN both have a 2-node hidden layer with 5 gridpoints (9648 parameters, see Table 3). Vertical green lines are the training data slices. All other data, including all magnitude values, are unseen during training. (B) Training loss comparison between LeanKAN and AddKAN. (C) Neural scaling results showing reduced MSE for LeanKAN at all tested network sizes.
The domain was discretized using $\Delta t=0.01$ and $\Delta x=0.05$, and solved using the Rodas5 stiff ODE integrator [6, 24]. We compare LeanKAN and AddKAN surrogate models, both of which contain two total layers and are of the structures introduced in Fig. 4(A). The task for which the KAN-based surrogate models are trained is learning the temporal term $\partial u/\partial t$ in Eq. 20 as a function of the entire current spatial solution profile, and then stepping forward to predict future states. Both surrogates have 402-dimensional input and output layers (doubled from the 201-dimensional grid to account for real and imaginary solution values). In a preliminary sparse case, we study for both AddKAN and LeanKAN a minimal two-node hidden layer, where all activations have just five gridpoints. We use the softsign normalizer in this case. The training data included eight evenly spaced snapshots between 0.1 s and 1.5 s of the real and imaginary solution values (see the first two entries in the top row of Fig. 6(A)), the training losses of which are shown in Fig. 6(B). Testing in Fig. 6(A) studies the two KAN-ODEs’ capabilities to reconstruct the entire spatiotemporal solution profile, as well as the derived magnitude fields. We see in this sparse representation using just two hidden nodes that the LeanKAN converges to nearly two orders of magnitude less training loss than the AddKAN, and that its convergence happens in much fewer epochs. While the training losses reveal the accuracy at the eight provided data slices, the complete reconstructed solution profiles demonstrate that the LeanKAN is also superior to the AddKAN in its ability to interpolate between the training data points and provide higher accuracy at unseen testing points.
While the LeanKAN studied here outperforms the identically structured AddKAN, both networks appear limited by the minimal two-layer hidden node. We additionally study the neural convergence of these two structures applied to the Schrödinger Equation in Fig. 6(C) and Table 3. Both KAN structures converge to strong loss metrics with just three nodes in the hidden layer, and then appear to plateau and become largely insensitive to further increases in size. In particular, LeanKAN outperforms AddKAN at all studied sizes. With three nodes appearing to be the optimal choice, further hyperparameter optimization is carried out, the results of which are shown in Table 4. In addition to the grid changes studied throughout the current work, we consider varying $n^{\text{mu}}$ values, the normalizer, and base activations. Interestingly, the optimal configuration based on these results appears to be that with $n^{\text{mu}}$=3, indicating multiplication only at the output layer (as well as the softsign normalizer, base activations on, and 10 gridpoints). The optimal $n^{\text{mu}}$ is likely problem-dependent and may require trial and error to determine. That being said, a broader look at the MSE results in Table 4 reveals that precise hyperparameter tuning may not be essential to the use of LeanKAN, where the best and worst results were roughly within an order of magnitude of variation (and the remaining results were fairly tightly clustered around the optimal value). We reiterate here that inclusion of the multiplication operator in LeanKAN adds a single hyperparameter only, compared to the two-dimensional optimization of $n^{a}$ and $k$ that would have been required when using MultKAN. In summary, we find in the current section that LeanKAN is also effective at PDE reconstruction, representing a notable efficiency improvement over previous KAN-ODE work that applied AddKANs only [13], while offering simple hyperparameter tuning.
Table 3: Neural convergence study for Schrödinger problem..
No. Params. Nodes Grid $n^{\text{mu}}$ LeanKAN MSE AddKAN MSE
9648 2 5 2 $1.24\times 10^{-5}$ $5.86\times 10^{-4}$
14472 3 5 2 $2.49\times 10^{-6}$ $5.59\times 10^{-6}$
26532 3 10 2 $1.81\times 10^{-6}$ $1.77\times 10^{-5}$
44220 5 10 3 $2.35\times 10^{-6}$ $5.37\times 10^{-6}$
Table 4: Hyperparameter study for Schrödinger problem.
Nodes Grid $n^{\text{mu}}$ Normalizer Base act. MSE
3 5 2 softsign On $2.49\times 10^{-6}$
3 10 2 softsign On $1.81\times 10^{-6}$
3 10 1 softsign On $5.82\times 10^{-6}$
3 10 3 softsign On $1.00\times 10^{-6}$
3 10 3 tanh On $2.71\times 10^{-6}$
3 10 3 softsign Off $1.15\times 10^{-5}$
5 Conclusions